OpenStudy (anonymous):

Find the derivative of y(theta)= theta+cos(^2)(3theta)

OpenStudy (fibonaccichick666):

so if you had an x instead of theta, do you think you could do it?

OpenStudy (anonymous):

i got 1-3sin(3theta) but i want to know why the answer is 1-3sin(6theta)

OpenStudy (danjs):

chain rule use

OpenStudy (fibonaccichick666):

ohk, well, I am not getting that answer because I donot know the double angle identities, but I can help ya through :)

OpenStudy (fibonaccichick666):

so we have \[y(\theta)= \theta+cos^2(3\theta)\]

OpenStudy (anonymous):

oh wow i dent know i had to use a double angle identity 0.o

OpenStudy (danjs):

dont need double angle

OpenStudy (fibonaccichick666):

first thing I want to do is substitute @DanJS at the end, that is how they simplified it

OpenStudy (danjs):

(-6)*sin(3x)*cos(3x)

OpenStudy (fibonaccichick666):

so @anarvizo, if we let \(u=3\theta\) can you tell me du=?

OpenStudy (anonymous):

so i'd get 1-3sin(3 theta) still ;/ unless I'm missing dropping the ^2 in from of sin

OpenStudy (danjs):

that is the derivative of cos^2(3x)

OpenStudy (anonymous):

okay lets see hang on

OpenStudy (danjs):

ic yeah you can simplify the answer with the double angle formula

OpenStudy (fibonaccichick666):

(I am showing you how to get that derivative)

OpenStudy (fibonaccichick666):

so what would du=?

OpenStudy (anonymous):

du would just = 3 wouldn't it?

OpenStudy (anonymous):

thats why i have the 3 in front of the 1-3sin..

OpenStudy (fibonaccichick666):

yep ok so that is important for later on when we apply the chain rule. Now, we need to rewrite a little for this to make sense

OpenStudy (danjs):

GL ... tag me if ya need anything

OpenStudy (fibonaccichick666):

\[y(\theta)= \theta+(cos(u))^2\]

OpenStudy (fibonaccichick666):

then I like to do a second substitution here

OpenStudy (fibonaccichick666):

let's say h=cosu what is dh=?

OpenStudy (anonymous):

do you always put the ^2 outside of a trig function when it says cos^2. sin^2, etc…?

OpenStudy (anonymous):

dh= -sin(u)

OpenStudy (solomonzelman):

There are 3 chains. \(\LARGE\color{black}{y=\theta+\cos^2(3\theta )}\) re-writing the function as, \(\LARGE\color{black}{y=\theta+[~\cos(3\theta )~]^2}\) (finding dy / d theta) \(\LARGE\color{black}{y=1+2[~\cos(3\theta )~]^{2-1}\times[-\sin(3\theta )]\times \theta}\) \(\LARGE\color{black}{y=1-2\theta\cos(3\theta )\sin(3\theta )}\)

OpenStudy (fibonaccichick666):

yes, it is just another way to write it.

OpenStudy (solomonzelman):

And you know that: \(\LARGE\color{black}{\sin(2a)=2\sin(a)\cos(a)}\)

OpenStudy (fibonaccichick666):

So now, we have \[y(\theta)= \theta+h^2\]

OpenStudy (anonymous):

oh writing it that way make much more sense lol now i see how theres multiple chain rules going on

OpenStudy (fibonaccichick666):

can you find the derivative of that?

OpenStudy (anonymous):

so now its 1+2h

OpenStudy (fibonaccichick666):

yea, that's why I think the substitution makes life easier. It's an extra step but it works

OpenStudy (fibonaccichick666):

ok so key thing here is that it y'=1+2h*dh*du

OpenStudy (fibonaccichick666):

that is your chain rule in work

OpenStudy (solomonzelman):

No, we don't need substitution. Do we?

OpenStudy (fibonaccichick666):

because the total derivative of h requires the previous derivatives too,(why we found them)

OpenStudy (solomonzelman):

ohh, I made an error.

OpenStudy (solomonzelman):

the derivative of 3theta is 3, and I wrote theta-:(

OpenStudy (solomonzelman):

sorry.

OpenStudy (fibonaccichick666):

so now, can you substitute everything in and tell me what you get @anarvizo

OpenStudy (fibonaccichick666):

@SolomonZelman , I like doing the substitutions because it makes you a lot less likely to mess up. I think They are a nice in between step that prevents little mistakes.

OpenStudy (solomonzelman):

I made a mistake not because of the method, but because of me:)

OpenStudy (anonymous):

okay so we have y'= 1+2(cos(theta))*-sin(theta)*3

OpenStudy (fibonaccichick666):

I wasn't implying you, sorry, I mean for everyone me included. I feel that way

OpenStudy (fibonaccichick666):

what was u again?

OpenStudy (anonymous):

theta? right?

OpenStudy (fibonaccichick666):

no u=3theta

OpenStudy (solomonzelman):

yes, that is correct, but one mistake. y'= 1+2(cos(\(\large\color{black}{ 3 }\)theta))*-sin(\(\large\color{black}{ 3 }\)theta)*3

OpenStudy (fibonaccichick666):

if you scroll up, you'll see it

OpenStudy (anonymous):

ohh duhhhh lol silly me

OpenStudy (solomonzelman):

mistakes happen to all people. you have just now wintessed this.

OpenStudy (fibonaccichick666):

lol it's ok, but so now that is right, I'll let Solomon explain the double angle simplification. Since he already did it. also wiNTessed? :)

OpenStudy (solomonzelman):

yeah, I am like dude... when I wrote it.

OpenStudy (anonymous):

how does that equal to 1-3sin(6theta) though? I have an extra cos unless theres an error or the answer document? hahaha and its fine :) human beings ar ethe best at making mistakes

OpenStudy (anonymous):

on*

OpenStudy (fibonaccichick666):

no no you are correct that answer is correct

OpenStudy (fibonaccichick666):

it's just one more step to make it pretty

OpenStudy (solomonzelman):

we are applying. \(\large\color{black}{ \sin(2\theta)=2\sin(\theta)\cos(\theta)}\)

OpenStudy (anonymous):

okay :)

OpenStudy (solomonzelman):

which in your case is, \(\large\color{black}{ \sin(6\theta)=2\sin(3\theta)\cos(3\theta)}\)

OpenStudy (anonymous):

ohh i see where you're going

OpenStudy (solomonzelman):

you can always prove this identity, it is just sin(a)cos(b)+sin(b)cos(a) where a and b are same.

OpenStudy (solomonzelman):

and after, you know that: sin(X)cos(X)+sin(X)cos(X)=2sin(X)cos(X).

OpenStudy (anonymous):

alrighty

OpenStudy (fibonaccichick666):

So it is just applying a known formula to make the answer pretty, both answers are equivalent though

OpenStudy (solomonzelman):

yeah, but the one with and angle of 6theta is shorter to write.

OpenStudy (fibonaccichick666):

true :) but I never memorized those formulas haha never needed to

OpenStudy (anonymous):

i had them memorized in precalc; not much anymore :/

OpenStudy (fibonaccichick666):

I never took pre-calc.... I skipped it

OpenStudy (solomonzelman):

I memorize everything that seems logical. I think double angles for sin and cos are good.

OpenStudy (solomonzelman):

smart..

OpenStudy (fibonaccichick666):

I just didn't use them enough to warrant memorizing them.

OpenStudy (anonymous):

i did it in high school :S I'm a freshman in college lol and yes very smart

OpenStudy (fibonaccichick666):

I did calc 1 and 2 in hs, we actually didn't have a pre-algebra class. just trig

OpenStudy (anonymous):

oh wow! AP calc was the only higher math course offered at my hs 0.o my teacher wasn't all that great though so here i am… lol. thanks for helping though!! all of you :)

OpenStudy (fibonaccichick666):

If you ever get confused or want to get ahead, what saved me in those was khanacademy.org and MITopencourseware. You can watch the lectures for the subject for free and get a better teacher. Khan is good if you are super confused, MIT is good for if you really understand the last section covered.

OpenStudy (fibonaccichick666):

np, and be warned college professors can be [extraordinarily]worse than hs, so try to prep for math classes. either watch the lectures or read the book if you understand it, that way during class you're not just like....uhhhmmm what's with all this greek? :)

OpenStudy (anonymous):

I've come to notice that about the professors D: lol thanks! haha quick question though; say i have to find where that slope =1… i'd just set the derivative equal to 1 and solve for theta..? I feel like its wrong though cause shouldn't there be multiple answers..?

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (solomonzelman):

I would think so:)

OpenStudy (solomonzelman):

knowing that the derivative is the slope.

OpenStudy (anonymous):

okay yeah that what made sense to me; thanks! wanted to make sure:)

OpenStudy (solomonzelman):

this is (similar to) what we do by a mean value theorem.

OpenStudy (fibonaccichick666):

I agree

OpenStudy (fibonaccichick666):

and MVT isn't covered until theoretical calc Solomon. I doubt they know it yet

OpenStudy (solomonzelman):

I covered it in Calc 1.

OpenStudy (solomonzelman):

Like I hate how they make separate course/semester for "abstract calc". it is now much different conceptually, besides that any problem would be very not-used to.

OpenStudy (solomonzelman):

i mean not much.

OpenStudy (anonymous):

the mean value theorem? we covered it :)

OpenStudy (solomonzelman):

lol, I also thought so:)

OpenStudy (fibonaccichick666):

really... hmm I don't remember doing it lol

OpenStudy (fibonaccichick666):

but then again, long time ago haha

OpenStudy (solomonzelman):

I haven't even heard of "theoretical calc"....

OpenStudy (fibonaccichick666):

.... I just passed it. It is awful. You have to prove everything. EVERYTHING from 1 variable calculus. In one semester

OpenStudy (fibonaccichick666):

So you prove limits, derivatives, integrals, but there is soooooo much info in between

OpenStudy (fibonaccichick666):

have you heard of \(\epsilon-\delta\) definition of a limit?